OpenStudy (firejay5):

Is y = 3x + 3 and y = 3x - 3 an inverse function? Explain why/how?

OpenStudy (anonymous):

are you asking if the two functions are inverses of each other?

OpenStudy (firejay5):

yes

OpenStudy (anonymous):

ok, so to find an inverse function, you solve a function for x, instead of y. then once you have something like x = _____. you then replace all the x's with y's and all the y's with x's so solve the first equation for x. then replace all the x's with y's and all the y's with x's. Is the resultant equation the same as the second equation. if so, then the two are inverse functions.

OpenStudy (firejay5):

are the two functions inverse

OpenStudy (anonymous):

try to find the inverse of the first function first, using the process I showed above.

OpenStudy (firejay5):

and then you'll tell me

OpenStudy (anonymous):

yes

OpenStudy (firejay5):

I already did and I couldn't tell

OpenStudy (anonymous):

y = 3x + 3 y-3 = 3x x = (y-3)/3 // change the x's for y's and the y's for x's y = (x-3)/3 so the inverse of the first function is not equal to the second function

OpenStudy (firejay5):

@saifoo.khan Is Tomo right?

OpenStudy (saifoo.khan):

Yes

OpenStudy (firejay5):

Can you explain why, because you're a better explainer

OpenStudy (firejay5):

Can you not explain it

OpenStudy (saifoo.khan):

Explaination is there int he video.

OpenStudy (firejay5):

I know I watched the video already didn't help

OpenStudy (saifoo.khan):

Actually we have to solve it this way: y = 3x + 3 Our goal is to make 'x' the subject first. After solving you should get: x = (y-3)/3 now simply change x as y and y as x. y = (x-3)/3 Hence that's our inverse. @tomo showed the same thing too.

OpenStudy (firejay5):

can you show it graphical

OpenStudy (saifoo.khan):

For graphing representation, we reflect the line to line y=x. See that video to understand it.

OpenStudy (firejay5):

Are you sure it's right? Answer is: y = 3x + 3 y = x - 3 / 3. @saifoo.khan

OpenStudy (firejay5):

@tomo what's the inverse of y = -3 is it x = -3

OpenStudy (firejay5):

@genius12 Is the inverse of y = -3, x = -3